OpenStudy (anonymous):

Given 27-x^3 what's f of f I put 27-27-x^3 but it was wrong

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you mean given \[f(x)=27-x^3\] what is \[f(f(x))\]??

OpenStudy (anonymous):

\[f(f(x))=f(27-x^3)=27-(27-x^3)^3\]